# Personalizing story generation by inferring author styles

### v1

@@ E.2 Effect of Persona on Similarity to Author Story
Table 9: Example showing the role of persona descriptions in determining similarity to the Author Story for a Reddit author. The Writing Sheet creatively blends fantasy and military strategy, making it more similar to the Author Story than the Average Author, while the Writing Sheet nP focuses primarily on direct conflict resolution.
Table 9 highlights the importance of persona descriptions in achieving similarity to the Author Story for a Reddit author. The Author Story creatively blends military strategy and fantasy, exemplified by phrases like “the weapons and armors of the enemy shifted by his will.” Similarly, the Writing Sheet integrates both elements, as seen in “It was as if the castle was alive, an ally in his solitary defense.” In contrast, the Average Author and Writing Sheet nP primarily focus on directly resolving the writing prompt through descriptions of battle. The Average Author incorporates minor fantasy elements, such as “amplified by some unseen force,” whereas the Writing Sheet nP adheres strictly to conventional military tactics. As a result, the Writing Sheet is preferred over the Average Author, which in turn is preferred over the Writing Sheet nP. This outcome demonstrates that following persona descriptions enhances alignment with the author’s style, as the Writing Sheet successfully captures their characteristic genre-blending approach of including elements of fantasy in the story. The broader thematic variety of Reddit, i.e., war-themed prompts, led to a more descriptive Author Writing Sheet capturing the author’s style of combining fantasy with the writing prompt, which helped construct a useful persona to enable better personalization using the Writing Sheet method.
E.3 Category-Wise Results for Similarity to Author Story
@@
(c) N.Magazine
(d) New Yorker
Figure 22: Win-rate proportion of our personalization methods against Average Author for AO3, Storium, N.Magazine, and New Yorker across four narrative categories for similarity to author story. D stands for Delta, S for Writing Summary, and WS for Writing Sheet.
Table 10: Win-rate percentages across narrative categories for similarity to the author story, averaged across sources. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
Category Delta Writing Sheet Writing Summary
Plot 24-37 32-50 34-45
Creativity 43-24 52-32 53-24
Development 32-39 37-39 40-43
Language Use 45-31 47-37 54-32
Table 10 shows that Delta, Writing Summary, and Writing Sheet outperform Average Author in Creativity (23% higher) and Language Use (15% higher) (Huot et al., 2024), but perform worse in Plot and slightly worse in Development (Tian et al., 2024; Xu et al., 2024). Creativity and Language Use are less dependent on the writing prompt, making them easier to transfer across prompts, whereas Plot and Development are more closely tied to specific prompts, making generalization difficult, especially without thematic overlap.
To better understand narrative category-wise personalization performance, Figure 22 shows the win-rates for sources - AO3, Storium, N.Magazine, and New Yorker similar to Figure 2 in the main paper. Overall, across all sources, we see that our proposed methods of Writing Sheet and Summary perform better than Delta across all four narrative categories. We see that Writing Summary generally performs better than Writing Sheet for sources like AO3 indicating that for the fanfiction domain explicitly contrasting the author-written story with the Average Story does not help much as the authors demonstrate limited stylistic variations among each other. In general, we observe that the win-rates of our personalization methods are higher than the Average Author for narrative categories of Creativity and Language Use as compared to Plot and Development confirming with prior results likely because Creativity and Language Use are less tied to the writing prompt than Plot and Development.
E.4 Traditional Metrics
Table 11: Comparison of methods across Lexical similarity, Diversity, and Style metrics.
@@ E.4 Traditional Metrics
Ground Truth 0.122 0.698 0.8882 1516
Oracle 0.0306 0.1577 0.1436 0.7132 0.7662 0.7712 1300
Average Author 0.0123 0.1455 0.1412 0.7146 0.7378 0.7289 1383
RAG 0.0118 0.1471 0.1432 0.7176 0.7393 0.7333 1326
Delta 0.0115 0.146 0.147 0.7182 0.7462 0.7388 1298
Writing Sheet 0.0101 0.1465 0.1452 0.7182 0.7735 0.7545 1344
Writing Summary 0.0109 0.1463 0.1474 0.7202 0.7589 0.7418 1353
Writing Sheet nP 0.0116 0.1475 0.1442 0.7168 0.7741 0.7563 1331
Writing Summary nP 0.0112 0.1461 0.1456 0.7176 0.768 0.7465 1313
Table 11 reports results using traditional metrics (Section D.2). Lexical Overlap and Diversity metrics yield similar values, as all methods use the same generation model, leading to overlapping lexical distributions and limiting these metrics’ ability to capture nuanced stylistic differences (Zheng et al., 2023; Xie et al., 2023). However, Writing Sheet nP consistently outperforms other methods, particularly in Style metrics for both similarity to Author History and Author Story, as measured by LUAR (Rivera-Soto et al., 2021). This improvement likely results from the Writing Sheet explicitly summarizing an author’s stylistic deviations from an Average Author, enhancing personalization. Homogenization scores are slightly worse than the Average Author, likely explained by the increased similarity among the generated stories for the same author.
@@ Appendix F LLama Results
Table 12: Percentage win-rate for Faithfulness to Author History for LLama 3.1 8B. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
AO3 34-59 30-59 54-40 51-34 45-46 51-40 50-36
Reddit 39-40 39-54 46-44 54-30 56-35 47-46 54-40
Storium 22-65 32-55 45-42 40-40 45-45 42-45 38-42
N.Magazine 14-64 21-71 29-43 57-43 57-29 29-64 21-57
New Yorker 12-80 20-73 40-60 53-33 53-27 40-47 47-40
Overall 23-62 28-62 43-46 51-36 51-36 42-48 42-43
Table 13: Percentage win-rate for Faithfulness to Writing History for LLama 3.1 70B. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
AO3 41-48 44-42 35-52 41-36 57-29 50-36 64-26
Reddit 42-46 28-65 40-44 53-37 53-35 53-37 67-21
Storium 30-57 52-32 38-50 35-40 40-38 25-60 48-35
N.Magazine 50-29 29-64 43-57 50-36 50-50 64-14 64-29
New Yorker 53-47 33-67 27-60 67-20 40-53 73-13 60-27
Overall 43-45 37-54 37-53 49-34 48-41 53-32 61-28
Table 14: Percentage win-rate for Similarity to Author Story for LLama 3.1 8B. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
AO3 66-26 31-51 52-34 48-38 38-50 52-34 51-32
Reddit 75-23 32-51 51-37 42-49 37-40 39-49 46-44
Storium 42-50 25-60 30-57 18-68 30-52 20-57 28-60
N.Magazine 43-50 36-43 43-57 21-50 36-64 43-57 36-43
New Yorker 27-60 07-80 13-60 20-67 13-67 0-93 20-73
Overall 51-42 26-57 38-49 30-54 31-55 31-58 36-50
Table 15: Percentage win-rate for Similarity to Author Story for LLama 3.1 70B. Each cell (‘X-Y’) shows ‘X’ as the method win-rate and ‘Y’ as the Average Author win-rate, with ties as the remainder.
Source Oracle RAG Delta Writing Sheet Writing Summary Writing Sheet nP Writing Summary nP
AO3 62-24 35-50 35-55 36-46 44-40 46-39 39-40
Reddit 74-21 37-49 40-42 47-39 49-37 42-44 44-33
Storium 50-42 30-57 32-52 18-57 38-52 28-62 42-40
N.Magazine 79-7 43-21 36-50 29-43 36-50 21-43 50-29
New Yorker 60-27 20-60 27-60 27-67 20-80 20-60 27-67
Overall 65-24 33-47 34-52 31-50 37-52 31-50 40-42
@@
Tables 12 and 13 show Faithfulness to Writing History for Llama 3.1 8B and 70B, while Tables 14 and 15 report similarity to the author story. Across both criteria, Oracle achieves higher win-rates with Llama 3.1 70B, indicating that larger models enhance instruction-following and personalization (Chung et al., 2024).
@@
For Llama 3.1 8B (Table 12), Writing Sheet and Writing Summary achieve the highest scores (9% higher than nP), demonstrating the benefit of persona descriptions. However, for Llama 3.1 70B (Table 13), Writing Sheet nP and Writing Summary nP outperform their persona-based counterparts by approximately 9%, suggesting that GPT-4o-generated personas provide no benefit to other models with stronger instruction-following capabilities (McKenzie et al., 2023). This is likely because persona descriptions of an author’s writing style are model-dependent and do not transfer well to other models with comparable performance (Shashidhar et al., 2024).
@@
For both Llama 3.1 8B (Table 14) and Llama 3.1 70B (Table 15), Writing Summary outperforms Writing Sheet by 5% and 9%, respectively (Shashidhar et al., 2024). This outcome likely stems from the construction process: the Author Writing Sheet contrasts an LLM-generated Average Story with the author-written story, making it more model-dependent. In contrast, the Author Writing Summary does not rely on the LLM-generated Average Author story, allowing for greater generalization across models.
@@ G.1 Experiment Design and Cost
We recruited three annotators via Upwork, compensating them at a rate of \$17 per hour. The annotators provided consent by signing a Google form that outlined the instructions of the task including the purpose of the research study. Each annotator evaluated 25 author stories and 75 story pairs, corresponding to the three personalization methods: Delta, Writing Sheet, and Writing Summary. Collectively, our annotators evaluated 45 unique author stories (135 story pairs), of which 15 author stories were annotated by all three annotators to assess inter-annotator agreement. The remaining 30 author stories were distributed uniquely, with each annotator evaluating 10 exclusive stories to increase annotation coverage (Song et al., 2024). The annotation process required a total of 12 hours and 30 minutes per annotator, averaging 30 minutes per author story. Each annotator received \$213 for their work, bringing the total annotation cost to \$639.
@@ G.2 Annotator Comments
"It’s terminal," the doctor told you as you, the one renowned as the world’s greatest hero, cradled your dying son, "sometimes, when the super genes mix, the outcome isn’t as expected." The room felt like it was closing in on me, the sterile smell of antiseptic failing to mask the sour stench of fear. My son, little Ethan, lay pale and fragile in my arms, his once vibrant eyes now dull and glassy. This was not how it was supposed to go. Heroes don’t have children who die. Heroes don’t face problems they can’t punch, fly away from, or blast with heat vision. “Isn’t there something we can do?” My voice cracked, and I hated how weak it sounded…….Gently, I laid him down, wrapping him in his favorite blanket. The sky was painted in hues of orange and pink, a beautiful contrast to the sorrow in my heart. I sat there, watching the sun dip below the horizon, and for the first time, I felt truly mortal. I realized that even heroes are not immune to the pain of loss, the fragility of life…….And as the night embraced the world, I knew that Ethan’s spirit would live on, a beacon of hope and courage in the heart of the world’s greatest hero.
Writing Sheet
@@
Table 17: Qualitative Example where Writing Sheet method is preferred over Average Author by all three annotators for similarity to the ground-truth author story for a Reddit author.
@@
Table 18: Qualitative Example where the Average Author method is preferred over Writing Sheet by all three annotators for similarity to the ground-truth author story for an AO3 author.
@@
Table 19: Qualitative Example where all three annotators select ‘Tie’ for similarity to the author story of a Storium author.
Writing Prompt
Ground-truth Author Story
Average Author
"When the word ’hospice’ is first spoken, it feels like the air itself changes, heavy with finality…It’s her mother, Marjorie, who the doctor is addressing, but the word reverberates through Clara’s own heart, a dull thud of inevitability." Marjorie stares at the doctor with a mix of resignation and defiance, her once-vibrant green eyes now clouded by the long battle with illness….. Her touch is light, almost ethereal, and Clara feels a sudden rush of tears she struggles to hold back. “Clara,” Marjorie says softly, her voice barely a whisper. “Do you remember the garden?” Clara nods, unable to speak. The garden had been Marjorie’s sanctuary, a riot of color and life where she had spent countless hours tending to roses, tulips, and wildflowers…… Marjorie’s decline is gradual but inevitable, and Clara feels the weight of each passing day. She continues to read to her mother, to sit by her side, and to hold her hand through the long nights. One crisp morning, as the first frost dusts the garden in a delicate shimmer, Marjorie slips away. Clara is by her side, her hand still entwined with her mother’s, and she feels a profound stillness settle over her…..
Writing Sheet
Table 20: Qualitative Example where all annotators select different winners for similarity to the author story of a New Yorker author.
@@ G.3 Qualitative Examples
We present qualitative examples from our human annotation process, covering various cases: instances where all annotators prefer Writing Sheet, all prefer Average Author, or select ‘Tie.’ Additionally, we include an example where annotators disagree, choosing different winners, highlighting the challenging and subjective nature of the annotation task.
All annotators prefer Writing Sheet:
Table 17 shows an example where all three annotators prefer the Writing Sheet method over the Average Author for similarity to the author story of a Reddit author. The annotators highlight several key aspects that make the Writing Sheet story more aligned with the ground-truth author story: (1) The Writing Sheet method incorporates humor and satire, using phrases such as “a tale of capes and cereal,” “invented breakfast hurricane,” and “World’s Greatest Hero Faces Greatest Challenge Yet: Parenting,” which mirror the lighthearted and satirical tone of the author story. (2) Unlike the Average Author story, which concludes with the son’s death, the Writing Sheet story reinterprets the prompt, portraying Titania as an advocate for families with special children. This aligns with the ground-truth author story’s approach, where the son engages with coding and the narrative avoids a sorrowful ending. (3) In terms of language style, the Writing Sheet story employs imaginative elements such as “The Hero’s Journey: A Tale of Capes and Cereal” and “The Avalon Gazette,” which parallel the ground-truth author story’s unconventional use of computer programming references, such as “Syntax Error: Unexpected Power.”
@@
Table 18 shows an example where all three annotators prefer the Average Author method over the Writing Sheet for similarity to the author story of an AO3 author. The primary reason cited by annotators is that the Writing Sheet method introduces a strong internal conflict for the king, portraying him as guilty or hesitant about sneaking into the kitchen for a late-night snack. His character is depicted as struggling with the decision, given his stature as a formidable ruler expected to lead battles. Additionally, the Writing Sheet creates a sharp contrast between the king and queen, with the queen depicted as joyful and mischievous, encouraging the king to abandon his guilt.
In contrast, both the ground-truth Author Story and the Average Author story emphasize the playful banter and affectionate dynamic between the king and queen, focusing on the shared experience of sneaking into the kitchen rather than introducing introspection or internal conflict. This shift in focus stems from the Writing Sheet being influenced by the author’s past writing history, where internal conflicts are a recurring theme. However, in this case, character development—being more writing-prompt-specific than broader categories like creativity or language use—did not transfer well. The author’s original story prioritizes external actions and lighthearted interactions rather than deep introspection, leading to a mismatch between the Writing Sheet’s personalized adaptation and the intended tone of the new prompt.
@@
Table 19 shows an example where all three annotators select ‘Tie’ for similarity to the author story of a Storium author. The primary reason is that both the Average Author story and the Writing Sheet story deviate significantly from the ground-truth author story in terms of plot. The ground-truth story adopts an open-ended approach to both character development and plot, where character relationships remain undefined, and the narrative focuses on anticipation rather than action.
In contrast, the Average Author story follows a conventional narrative arc, featuring fast-paced dialogue and character interactions that build tension and lead to resolution. Similarly, the Writing Sheet story also resolves the narrative but differs from the Average Author method by emphasizing past missions and their consequences, aligning with the author’s tendency to reference character backstories. However, both methods fail to capture the open-ended nature of the original story. This misalignment in plot structure leads all three annotators to select ‘Tie’ for this example.
@@
The annotator who selected Writing Sheet as the winner states, ‘‘Although Average Author and the Author Story depict relationships between close family members, Writing Sheet is more similar to the Author Story in Plot, Creativity, and Development by exploring the meaning and emotional responses to the term hospice.’’ This annotator highlights the presence of subtext in the Writing Sheet story and its attempts to mimic the author’s cyclic style—“Each door held a story, a life in its final chapter… she would ask the same question…” and “Clara knew that work was not about erasing the pain or the grief, but about finding moments of grace within it…”.
The annotator who selected Average Author as the winner states, ‘‘Average Author aligns more closely with the Author’s Story in terms of Plot and Language Use. This is because both stories center on someone taking care of a family member in their own home, and growing from a lack of peace into a sense of peace, which is expressed through the choice of language. Writing Sheet also indicates a development from lack of peace to acceptance, but the familial tie is what pushes Average Author over the top.’’

### v2

@@ E.2.3 Effect of Persona on Similarity to Author Story
Table 15: Example showing the role of persona descriptions in determining similarity to the Author Story for a Reddit author. The Sheet creatively blends fantasy and military strategy, making it more similar to the Author Story than the Average Author, while the Sheet nP focuses primarily on direct conflict resolution.
Table 15 highlights the importance of persona descriptions in achieving similarity to the Author Story for a Reddit author. The Author Story creatively blends military strategy and fantasy, exemplified by phrases like “the weapons and armors of the enemy shifted by his will.” Similarly, the Sheet integrates both elements, as seen in “It was as if the castle was alive, an ally in his solitary defense.” In contrast, the Average Author and Sheet nP primarily focus on directly resolving the writing prompt through descriptions of battle. The Average Author incorporates minor fantasy elements, such as “amplified by some unseen force,” whereas the Sheet nP adheres strictly to conventional military tactics. As a result, the Sheet is preferred over the Average Author, which in turn is preferred over the Sheet nP. This outcome demonstrates that following persona descriptions enhances alignment with the author’s style, as the Sheet successfully captures their characteristic genre-blending approach of including elements of fantasy in the story. The broader thematic variety of Reddit, i.e., war-themed prompts, led to a more descriptive Author Writing Sheet capturing the author’s style of combining fantasy with the writing prompt, which helped construct a useful persona to enable better personalization using the Sheet method.
E.2.4 Category-Wise Results for Similarity to Author Story
@@
(c) N.Mag
(d) N.York
Figure 22: Win-rates proportions across narrative categories for Similarity to Author Story for AO3, Storium, N.Mag, and N.York across four narrative categories for similarity to author story.O: Oracle, D: Delta, Sh: Sheet, Su: Summ.
Figure 22 shows win-rate proportions across narrative categories for Similarity to Author Story on AO3, Storium, N.Mag, and N.York. Similar to Reddit, Creativity and Language Use show higher personalization performance than Plot and Development, because the latter are more dependent on the specific writing prompt.
E.3 Traditional Metrics
@@
Ground Truth 0.1222 0.7015 0.7402 1074
Oracle 0.0311 0.1577 0.1427 0.7167 0.6385 0.7707 930
Average Author 0.0122 0.1448 0.1415 0.7190 0.6149 0.7306 985
RAG 0.0117 0.1463 0.1428 0.7212 0.6161 0.7359 947
Delta 0.0114 0.1450 0.1462 0.7217 0.6218 0.7391 929
Sheet 0.0102 0.1454 0.1440 0.7215 0.6446 0.7532 959
Sheet-nP 0.0114 0.1464 0.1433 0.7208 0.6451 0.7544 950
Summ 0.0108 0.1451 0.1460 0.7237 0.6324 0.7432 965
Summ-nP 0.0111 0.1450 0.1445 0.7222 0.6400 0.7459 939
Table 16: Comparison of methods across Lexical similarity, Diversity, and Style metrics for GPT-4o generated stories.
Table 16 reports results using traditional metrics (Section D.2). Lexical Overlap and Diversity metrics yield similar values, as all methods use the same generation model, leading to overlapping lexical distributions and limiting these metrics’ ability to capture nuanced stylistic differences (Zheng et al., 2023; Xie et al., 2023). However, Sheet-nP consistently outperforms other methods, particularly in Style metrics for both similarity to Author History and Author Story, as measured by LUAR (Rivera-Soto et al., 2021). This improvement likely results from the Sheet explicitly summarizing an author’s stylistic deviations from an Average Author, enhancing personalization. Homogenization scores are slightly worse than the Average Author, likely explained by the increased similarity among the generated stories for the same author.
@@ Appendix F LLama Results
F.1 Faithfulness to Author History
Method Reddit AO3 Storium N.Mag N.York All
RAG 39 30 32 21 20 28
Delta 46 54 45 29 40 43
Sheet 54 51 40 57 53 51
Sheet-nP 47 51 42 29 40 42
Summ 56 45 45 57 53 51
Summ-nP 54 50 38 21 47 42
Table 17: Win-rates (%) for Faithfulness to Writing History vs. Average Author baseline for LLaMA 3.1 8B evaluated by GPT-4o.
Method Reddit AO3 Storium N.Mag N.York All
RAG 28 44 52 29 33 37
Delta 40 35 38 43 27 37
Sheet 53 41 35 50 67 49
Sheet-nP 53 50 25 64 73 53
Summ 53 57 40 50 40 48
Summ-nP 67 64 48 64 60 61
Table 18: Win-rates (%) for Faithfulness to Writing History vs. Average Author baseline for LLaMA 3.1 70B evaluated by GPT-4o.
@@ F.1 Faithfulness to Author History
Tables 17 and 18 show Faithfulness to Writing History for Llama 3.1 8B and 70B, while Tables 19 and 20 report similarity to the author story. Across both criteria, Oracle achieves higher win-rates with Llama 3.1 70B, indicating that larger models enhance instruction-following and personalization (Chung et al., 2024).
@@
For Llama 3.1 8B (Table 17), Sheet and Summ achieve the highest scores (9% higher than nP), demonstrating the benefit of persona descriptions. However, for Llama 3.1 70B (Table 18), Sheet nP and Summ nP outperform their persona-based counterparts by approximately 9%, suggesting that GPT-4o-generated personas provide no benefit to other models with stronger instruction-following capabilities (McKenzie et al., 2023). This is likely because persona descriptions of an author’s writing style are model-dependent and do not transfer well to other models with comparable performance (Shashidhar et al., 2024).
F.2 Similarity to Author Story
Method Reddit AO3 Storium N.Mag N.York All
Oracle 75 66 42 43 27 51
RAG 32 31 25 36 7 26
Delta 51 52 30 43 13 38
Sheet 42 48 18 21 20 30
Sheet-nP 39 52 20 43 0 31
Summ 37 38 30 36 13 31
Summ-nP 46 51 28 36 20 36
Table 19: Win-rates (%) for Similarity to Author Story vs. Average Author baseline for LLaMA 3.1 8B evaluated by GPT-4o.
Method Reddit AO3 Storium N.Mag N.York All
Oracle 74 62 50 79 60 65
RAG 37 35 30 43 20 33
Delta 40 35 32 36 27 34
Sheet 47 36 18 29 27 31
Sheet-nP 42 46 28 21 20 31
Summ 49 44 38 36 20 37
Summ-nP 44 39 42 50 27 40
Table 20: Win-rates (%) for Similarity to Author Story vs. Average Author baseline for LLaMA 3.1 70B evaluated by GPT-4o.
Method Sheet Sheet-ft Summ Summ-ft
Reddit (%) 42 62 37 60
Table 21: Win-rates (%) for Similarity to Author Story with and without fine-tuning for LLaMA 3.1 8B evaluated by GPT-4o for the Reddit subset of our dataset. Fine-tuned (ft) models show significant improvements over their counterparts with just prompting.
@@ F.2 Similarity to Author Story
For both Llama 3.1 8B (Table 19) and Llama 3.1 70B (Table 20), Summ outperforms Sheet by 5% and 9%, respectively (Shashidhar et al., 2024). This outcome likely stems from the construction process: the Author Writing Sheet contrasts an LLM-generated Average Story with the author-written story, making it more model-dependent. In contrast, the Author Writing Summary does not rely on the LLM-generated Average Author story, allowing for greater generalization across models.
Fine-tuning on the Author Writing Sheet/ Summary boosts performance for LLaMA 3.1 8B:
We experiment with fine-tuning the LLaMA 3.1 8B model on the Author Writing Sheet and Summary for the Reddit subset by prepending the Author Writing Sheet/ Summary to the writing prompt during story generation, in addition to prompting alone (see Table 21). We compare the fine-tuned method against an equivalent Average Author baseline, fine-tuned on the writing prompt alone without the Author Writing Sheet/ Summary. Our method with Sheet achieves a 62% win-rate against the baseline, compared to 42% (‘Sheet’ row of Table 19) for the prompting-only approach, yielding a 20% improvement. Similarly, for Summ, we achieve a 60% win-rate against the baseline, compared to 37% (‘Summ’ row of Table 19) for the prompting-only approach, yielding a 23% improvement. These results highlight the potential of Author Writing Sheet/ Summary generated by a different model (GPT-4o) to effectively guide another model (LLaMA 3.1 8B) when fine-tuned rather than just prompted.
@@ G.1 Experiment Design and Cost
We recruited three annotators via Upwork, compensating them at a rate of \$17 per hour. The annotators provided consent by signing a Google form that outlined the instructions of the task including the purpose of the research study. Each annotator evaluated 25 author stories and 75 story pairs, corresponding to the three personalization methods: Delta, Sheet, and Summ. Collectively, our annotators evaluated 45 unique author stories (135 story pairs), of which 15 author stories were annotated by all three annotators to assess inter-annotator agreement. The remaining 30 author stories were distributed uniquely, with each annotator evaluating 10 exclusive stories to increase annotation coverage (Song et al., 2024). The annotation process required a total of 12 hours and 30 minutes per annotator, averaging 30 minutes per author story. Each annotator received \$213 for their work, bringing the total annotation cost to \$639.
@@ G.2 Annotator Comments
"It’s terminal," the doctor told you as you, the one renowned as the world’s greatest hero, cradled your dying son, "sometimes, when the super genes mix, the outcome isn’t as expected." The room felt like it was closing in on me, the sterile smell of antiseptic failing to mask the sour stench of fear. My son, little Ethan, lay pale and fragile in my arms, his once vibrant eyes now dull and glassy. This was not how it was supposed to go. Heroes don’t have children who die. Heroes don’t face problems they can’t punch, fly away from, or blast with heat vision. “Isn’t there something we can do?” My voice cracked, and I hated how weak it sounded…….Gently, I laid him down, wrapping him in his favorite blanket. The sky was painted in hues of orange and pink, a beautiful contrast to the sorrow in my heart.
Sheet
@@
Table 23: Qualitative example where the Sheet method is preferred over Average Author by all annotators for a Reddit author. Sheet captures the author’s creative use of humor, satire, and unconventional programming jargon, which the Average Author misses.
Writing Prompt
Ground-truth Author Story
Average Author
"When the word ’hospice’ is first spoken, it feels like the air itself changes, heavy with finality…It’s her mother, Marjorie, who the doctor is addressing, but the word reverberates through Clara’s own heart, a dull thud of inevitability." Marjorie stares at the doctor with a mix of resignation and defiance, her once-vibrant green eyes now clouded by the long battle with illness….. Her touch is light, almost ethereal, and Clara feels a sudden rush of tears she struggles to hold back. “Clara,” Marjorie says softly, her voice barely a whisper. “Do you remember the garden?” Clara nods, unable to speak. The garden had been Marjorie’s sanctuary, a riot of color and life where she had spent countless hours tending to roses, tulips, and wildflowers…… Marjorie’s decline is gradual but inevitable, and Clara feels the weight of each passing day. She continues to read to her mother, to sit by her side, and to hold her hand through the long nights.
Sheet
Table 24: Qualitative Example where annotators select different winners for a New Yorker author. The Author Story uses metaphor and subtext to reframe hospice as both a physical and emotional space. While Sheet captures these elements through cyclic structure and introspective tone, the subjective nature of subtext leads annotators to different interpretations of similarity.
@@
Table 25: Example where all annotators prefer the Average Author over Sheet for an AO3 author. The Average Author aligns better with the story’s playful tone and king–queen dynamic, aided by LLM familiarity with Lord of the Rings fanfiction tropes. Sheet adds internal conflict that doesn’t fit this prompt.
@@
Table 26: Qualitative Example where all annotators select ‘Tie’ for a Storium author. While the Author Story uses an open-ended narrative with unresolved character dynamics, both the Average Author and Sheet introduce resolution and tension-driven plots, missing the author’s intended tone.
@@ G.3 Qualitative Examples
We present qualitative examples from our human annotation process, covering various cases: instances where all annotators prefer Sheet, all prefer Average Author, or select ‘Tie.’ Additionally, we include an example where annotators disagree, choosing different winners, highlighting the challenging and subjective nature of the annotation task.
All annotators prefer Sheet:
Table 23 shows an example where all three annotators prefer the Sheet method over the Average Author for similarity to the author story of a Reddit author. The annotators highlight several key aspects that make the Sheet story more aligned with the ground-truth author story: (1) The Sheet method incorporates humor and satire, using phrases such as “a tale of capes and cereal,” “invented breakfast hurricane,” and “World’s Greatest Hero Faces Greatest Challenge Yet: Parenting,” which mirror the lighthearted and satirical tone of the author story. (2) Unlike the Average Author story, which concludes with the son’s death, the Sheet story reinterprets the prompt, portraying Titania as an advocate for families with special children. This aligns with the ground-truth author story’s approach, where the son engages with coding and the narrative avoids a sorrowful ending. (3) In terms of language style, the Sheet story employs imaginative elements such as “The Hero’s Journey: A Tale of Capes and Cereal” and “The Avalon Gazette,” which parallel the ground-truth author story’s unconventional use of computer programming references, such as “Syntax Error: Unexpected Power.”
@@
Table 25 shows an example where all three annotators prefer the Average Author method over the Sheet for similarity to the author story of an AO3 author. The primary reason cited by annotators is that the Sheet method introduces a strong internal conflict for the king, portraying him as guilty or hesitant about sneaking into the kitchen for a late-night snack. His character is depicted as struggling with the decision, given his stature as a formidable ruler expected to lead battles. Additionally, the Sheet creates a sharp contrast between the king and queen, with the queen depicted as joyful and mischievous, encouraging the king to abandon his guilt.
In contrast, both the ground-truth Author Story and the Average Author story emphasize the playful banter and affectionate dynamic between the king and queen, focusing on the shared experience of sneaking into the kitchen rather than introducing introspection or internal conflict. This shift in focus stems from the Sheet being influenced by the author’s past writing history, where internal conflicts are a recurring theme. However, in this case, character development being more writing-prompt-specific than broader categories like creativity or language use did not transfer well. The author’s original story prioritizes external actions and lighthearted interactions rather than deep introspection, leading to a mismatch between the Sheet’s personalized adaptation and the intended tone of the new prompt.
@@
Table 26 shows an example where all three annotators select ‘Tie’ for similarity to the author story of a Storium author. The primary reason is that both the Average Author story and the Sheet story deviate significantly from the ground-truth author story in terms of plot. The ground-truth story adopts an open-ended approach to both character development and plot, where character relationships remain undefined, and the narrative focuses on anticipation rather than action.
In contrast, the Average Author story follows a conventional narrative arc, featuring fast-paced dialogue and character interactions that build tension and lead to resolution. Similarly, the Sheet story also resolves the narrative but differs from the Average Author method by emphasizing past missions and their consequences, aligning with the author’s tendency to reference character backstories. However, both methods fail to capture the open-ended nature of the original story. This misalignment in plot structure leads all three annotators to select ‘Tie’ for this example.
@@
The annotator who selected Sheet as the winner states, ‘‘Although Average Author and the Author Story depict relationships between close family members, Sheet is more similar to the Author Story in Plot, Creativity, and Development by exploring the meaning and emotional responses to the term hospice.’’ This annotator highlights the presence of subtext in the Sheet story and its attempts to mimic the author’s cyclic style—“Each door held a story, a life in its final chapter… she would ask the same question…” and “Clara knew that work was not about erasing the pain or the grief, but about finding moments of grace within it…”.
The annotator who selected Average Author as the winner states, ‘‘Average Author aligns more closely with the Author’s Story in terms of Plot and Language Use. This is because both stories center on someone taking care of a family member in their own home, and growing from a lack of peace into a sense of peace, which is expressed through the choice of language. Sheet also indicates a development from lack of peace to acceptance, but the familial tie is what pushes Average Author over the top.’’